OpenStudy (aaronandyson):

A positive number is divided into two parts such that the sum of the square of the two parts is 20.The square of the larger part is 8 times the smaller part.Taking x as the smaller part of the two parts,find the number @perl @mathslover

OpenStudy (anonymous):

The two parts' square sums to 20. The only squares that some to 20 are 4 and 16. The smaller part (x) is therefore 2. This is confirmed by the fact that 2=16/8. So the square of the larger part is 8 times x. Since the number was divided to get the two parts we must multiply to find the number. 2 x 4=8 so the number is 8.

OpenStudy (aaronandyson):

???

OpenStudy (perl):

Let x = smaller number, y = bigger number sum of the squares of the numbers is 20, bigger number squared is 8 times smaller x^2 + y^2 = 20 y^2 = 8*x

OpenStudy (perl):

by substitution x^2 + 8x = 20 x^2 + 8x - 20 = 0 ( x - 2) ( x + 10) = 0 x = 2, -10 remember x is positive, so x is 2

OpenStudy (aaronandyson):

What?

OpenStudy (aaronandyson):

@perl

OpenStudy (alekos):

He's given you the answer.

OpenStudy (aaronandyson):

The book says it 3 and 17 :S

OpenStudy (alekos):

are you sure that s the right Answer? Have you got a screen shot of the question?

OpenStudy (mathmate):

@AaronAndyson Perhaps you want to check if the answer corresponds to the right question.

OpenStudy (perl):

I'm not sure how your book got that answer. You can verify my work here, they did the same exact thing http://www.algebra.com/algebra/homework/word/numbers/Numbers_Word_Problems.faq.question.887984.html

OpenStudy (alekos):

Looks like the cat's got your tongue

OpenStudy (aaronandyson):

Yes the question number correspond with the answer number

OpenStudy (alekos):

Well it must be a typo

OpenStudy (mathmate):

@AaronAndyson Thank you for checking. I believe there is a mistake in the answer, or the wrong page, or whatever. The question says: "the sum of the square of the two parts is 20" The square of 17 is 289, already exceeds 20, while the \(sum\) of the two parts is 20. So the answer cannot correspond to the posted question...unless there is a typo in the question itself. I think your teacher is the only one who can sort this out, assuming there is no typo on your part.

OpenStudy (aaronandyson):

Nevertheless,Can you explain my new question in DETAIL?